OpenStudy (lukecrayonz):

Complex Numbers in Trigonometric Form and DeMoivre's Theorem for Roots Represent the complex number graphically, and find the trigonometric form of the number.

OpenStudy (lukecrayonz):

-2-2i

OpenStudy (lukecrayonz):

So far: -2-2i, r=sqrt(4+4)=sqrt(8)=2sqrt(2) tan theta=-2/-2=1=7pi/4?

OpenStudy (lukecrayonz):

Or no because its not negative one..

OpenStudy (anonymous):

you need \[r=\sqrt{(-2)^2+(-2)^2}=\sqrt{8}=2\sqrt{2}\]

OpenStudy (lukecrayonz):

Which i've done :)

OpenStudy (anonymous):

you also need \[\theta =\frac{5\pi}{4}\] although of course it is not uniqu

OpenStudy (lukecrayonz):

Then i get tan theta= -2/-2=1. How did you get 5 pi/4?

OpenStudy (anonymous):



OpenStudy (lukecrayonz):

5pi/4=225 degrees, which I understand, but I'm not sure where you actually got the 5pi/4 from from tan=1.

OpenStudy (anonymous):

you just have to look and see where you are in the coordinate plane in this case you have both x and y negative, so you are to the left and down. that is how you find the angle you know \[\tan(\theta)=1\] but you have to look to see what quadrant you are in to find \[\theta\]

OpenStudy (lukecrayonz):

Is it tan theta=1 because -2sqrt(2)/-2sqrt(2)=1? Since tan=y/x?

OpenStudy (anonymous):

this has nothing to do with converting from degrees to radians. it is the measure of the angle you are looking for,and it is not unique

OpenStudy (lukecrayonz):

I feel like I'm either going way too into it, or not enough. But I was just stating the 225 degrees because of your drawing.

OpenStudy (anonymous):

\[a+bi=r\left(\cos(\theta)+i\sin(\theta)\right)\] \[r=\sqrt{a^2+b^2}\] \[\tan(\theta)=\frac{b}{a}\]

OpenStudy (anonymous):

if you are working in radians (aka numbers) forget degrees.

OpenStudy (anonymous):



OpenStudy (lukecrayonz):

So if tan theta=b/a, and b=4 and a=4, arctan(1)=theta?

OpenStudy (anonymous):

that represents and angle of \[\frac{5\pi}{4}\]

OpenStudy (anonymous):

wait slow

OpenStudy (lukecrayonz):

I'm using a different chapter of knowledge for this because my online textbook is reffering to a chapter I was not in the class for.

OpenStudy (anonymous):

\[\tan(\theta)=\frac{b}{a}\] but that does not mean \[\arctan(\frac{b}{a})=\theta\]!!

OpenStudy (lukecrayonz):

Well it works out in this case...

OpenStudy (anonymous):

no it does not

OpenStudy (lukecrayonz):

Here's the method to my madness, and we are getting the same answer:

OpenStudy (anonymous):

arctan has range from \[-\frac{\pi}{2}\] to \[\frac{\pi}{2}\] so it will only work if you are in quadrant I or IV

OpenStudy (lukecrayonz):

tan(theta)=b/a b=4 a=4 arctan(1)=theta 45 degrees=theta ASTC, 45+180=225 degrees.

OpenStudy (anonymous):

if you are in quadrant II or III you have to figure out what the angle is. you can use arctan, but if you are in the wrong quadrant you have to adjust

OpenStudy (anonymous):

acually \[a=-2, b=-2\] but you are right \[\frac{b}{a}=1\] and you are also rigth that you need to add \[\pi\] to \[\frac{\pi}{4}\] to get the right angle